# NuwaTS: a Foundation Model Mending Every Incomplete Time Series

### v2

@@ Abstract
Time series imputation plays a crucial role in various real-world systems and has been extensively explored. Models for time series imputation often require specialization, necessitating distinct designs for different domains and missing patterns. In this study, we introduce NuwaTS, a framework to repurpose Pre-trained Language Model (PLM) for general time series imputation. Once trained, this model can be applied to imputation tasks on incomplete time series from any domain with any missing patterns. We begin by devising specific embeddings for each sub-series patch of the incomplete time series. These embeddings encapsulate information about the patch itself, the missing data patterns within the patch, and the patch’s statistical characteristics. To enhance the model’s adaptability to different missing patterns, we propose a contrastive learning approach to make representations of the same patch more similar across different missing patterns. By combining this contrastive loss with the missing data imputation task, we train PLMs to obtain a one-for-all imputation model. Furthermore, we utilize a plug-and-play layer-wise fine-tuning approach to train domain-specific models. Experimental results demonstrate that leveraging a dataset of over seventeen million time series from diverse domains, we obtain a one-for-all imputation model which outperforms existing domain-specific models across various datasets and missing patterns. Additionally, we find that NuwaTS can be generalized to other time series tasks such as forecasting.
@@ 1 Introduction
Time series data are pervasive across numerous fields, including transportation [16], economics, healthcare [34], and meteorology [2]. However, real-world time series data often suffer from missing values. Incomplete data complicate various time series applications such as forecasting and classification, ultimately degrading the quality of data analysis [24]. This makes time series imputation especially critical. Traditionally, time series imputation methods have leaned heavily on statistical techniques like mean imputation and interpolation [36]. Yet, with the rise of deep learning, there has been a notable shift towards deep learning models for imputation tasks [6; 8]. Despite significant advancements in deep learning-based imputation methods, they often encounter limitations as they are generally trained and utilized within a single domain-specific dataset. For instance, a model trained on time series data from the transportation sector may not be effective for imputing missing data in other domains such as weather and electricity (Figure 1(a)).
In contrast, foundation models from the fields of Computer Vision (CV) [15] and Natural Language Processing (NLP) [4; 29], pre-trained on vast and diverse datasets from various domains, have shown remarkable versatility. These models can be adapted to a wide array of applications across different domains through methods such as fine-tuning [10; 45] or prompting [11; 15]. Inspired by these achievements, we aim to develop a versatile one-for-all model that can impute incomplete time series data from any domain and accommodate any pattern of missing data. (Figure 1(b)).
((a)) Specialized models are separately trained on time series domains with specific missing data patterns and are only applicable to their specialized domains and missing data patterns.
((b)) One-for-all models (Ours): learned from relatively complete time series across various domains. Once trained, they can be applied to impute any incomplete time series from any domain.
Figure 1: Comparison between one-for-all models and specialized models for missing data imputation on time series.
However, effectively learning a one-for-all imputation model for diverse domains is technically challenging. The model must possess strong adaptability to various domains and missing data patterns. Additionally, in some applications, it needs the capability to quickly specialize to a specific domain with few-shot learning while retaining its generalizability. To tackle these challenges, we have developed NuwaTS, a one-for-all model for incomplete time series, distinguished by several strategies: First, recognizing that patches based on sub-series offer richer semantic information than individual time points, we have designed specific embeddings for each sub-series patch. These tokens include information for the patch itself, the missing data patterns within the patch, and the patch’s statistical information, serving as inputs to the PLMs. Second, we introduced a contrastive learning module that encourages the model to produce more similar representations of the same patch under varying missing data patterns. The contrastive loss and the reconstruction loss together constitute the loss function we use to train the PLM parameters. Third, when a specialized model tailored to a specific domain is required, we have designed a domain-specific prefix embedding and developed a plug-and-play fine-tuning mechanism. This mechanism introduces modules that insert well-designed continuous prompts into each layer of the frozen pre-trained one-for-all model without altering any of its weights. Once these modules are removed, the model reverts to its original state as a versatile, one-for-all model.
Our contributions are as follows:
@@
    We introduce a novel solution, NuwaTS, capable of performing missing data imputation tasks on any incomplete time series unseen during the training phase. To the best of our knowledge, this represents the pioneering attempt in this field.
@@
    We present a “plug-and-play” fine-tuning technique that seamlessly transforms a one-for-all model into a domain-specific model with minimal data and computational resources, all without modifying the model’s original weights.
@@
    The one-for-all NuwaTS consistently outperforms domain-specific state-of-the-art methods in imputation tasks across nearly all missing rates. Moreover, fine-tuned NuwaTS can be extended to time series forecasting, where its forecasting results are comparable to or even better than existing domain-specific time series forecasting models.
@@ 2.1 Incomplete Time Series Imputation
Many time series imputation models are tailored to specific missing data patterns and domains, such as randomly missing traffic time series. For instance, matrix factorization models are designed for multivariate time series imputation [40]. For a dataset from a specific domain with a particular missing rate, it often requires optimizing a specialized model using a low-rank prior as the optimization target for imputation. Recent advancements in deep learning techniques have shown promising results in addressing missing data imputation. Generative models such as Generative Adversarial Networks (GANs)[23; 33] and diffusion models are used to learn the underlying distribution of the incomplete time series. Several architectures, such as Recurrent Neural Networks (RNNs)[6; 25; 22; 32] and attention mechanisms [8], are proposed to capture temporal dependencies within incomplete time series. While achieving good performance on narrow domains and missing data patterns, these models lack versatility and generalizability to diverse domains and missing data patterns.
@@ 2.2 Foundation Model for Time Series
In recent years, foundational models for time series analysis have made notable strides, primarily leveraging pre-trained backbones from NLP and aligning modalities to extend their reasoning capabilities to time series tasks. For example, Gruver et al. [9] discovered that by encoding time series as strings of numerical digits, LLMs can perform time series forecasting with zero-shot capabilities. GPT4TS [44] trains time series models using pre-trained GPT weights. UniTime [18] integrates domain-specific instructions into LLMs, enhancing their generalization across domains. TIME-LLM [12] employs LLMs’ reasoning by framing time series data and statistical information in textual prompts, aligning time patches with word embeddings via cross-attention mechanisms for improved zero-shot learning. Autotimes [21] utilizes precomputed text embedding as positional embeddings and an autoregressive approach for long-term forecasting. TEST [31] uses text-prototype-aligned embeddings to enhance LLMs’ reasoning in time series data. $S^{2}$-LLM [28] applies seasonal-trend decomposition to time series and uses semantic space-informed prompting to retrieve appropriate prompts from word token embeddings as prefixes. LLM4TS [3] adapts LLMs for time series representation learning by directly converting individual patches into time series sequences. Chronos [1] trains language models from scratch on a large collection of time series data, using scaling and quantization for tokenization. This model demonstrates strong capabilities in zero-shot probabilistic forecasting. These models primarily focus on time series forecasting and do not specifically address missing data issues or the design of embeddings for missing data patterns.
@@ 3.1 Preliminaries and Problem Statement
$m_{t}=\begin{cases}1&\text{if }x_{t}\text{ is observed}\\ 0&\text{otherwise}\end{cases}$ (1)
Suppose we have acquired $N$ relatively complete time series datasets $\mathcal{D}=\{\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{N}\}$ from a diverse set of domains $\{\mathcal{S}^{1},\mathcal{S}^{2},\ldots,\mathcal{S}^{K}\}$, where $\mathbf{x}^{n}\in\mathbb{R}^{T_{n}}$ represents the $n^{th}$ time series with length $T_{n}$. Here, a “relatively complete time series” implies that $\sum{{m}_{t}^{n}}\ll T_{n}$, where $\mathbf{m}^{n}$ denotes the missing data pattern in the $n$-th series. Our objective is to utilize $\mathcal{D}$ to train an imputation model, denoted as $f_{\mathbf{\Phi}}$, characterized by parameters $\mathbf{\Phi}$. For any incomplete time series $\hat{\mathbf{x}}\in\mathbb{R}^{\hat{T}}$ accompanied by any missing pattern $\hat{\mathbf{m}}\in\mathbb{R}^{\hat{T}}$ from any domain $\hat{\mathcal{S}}$, the model $f_{\mathbf{\Phi}}$ can impute the missing values in $\hat{\mathbf{x}}$ as accurately as possible.
@@ 3.2 Model Architecture
3.2.1 Instance Normalization & Patching
Time series from different domains may exhibit variations in magnitude and distribution. To mitigate these discrepancies, we apply reversible instance normalization [14] to each variable $\mathbf{x}_{i}\in\mathbb{R}^{L}$ (where $\mathbf{x}_{i}$ is a segment of length $L$ randomly extracted from a time series in the dataset $\mathcal{D}$) before inputting it into the model. For the missing values, we assign them a value of 00. Recognizing that individual time points lack sufficient semantic meaning and hinder the model’s ability to recognize domain information, we utilize patching techniques. We define the patch length as $P$ and the non-overlapping region between two consecutive patches as $S$. The patching process then generates a sequence of patches $\mathbf{X}_{i,(p)}\in\mathbb{R}^{P\times N}$, where $N$ is the number of patches, calculated as $N=\left\lfloor\frac{L-P}{S}\right\rfloor+2$. Additionally, we pad the sequence with $S$ repeated values of the last data point $\mathbf{x}_{i}$ before initiating the patching process. We then use a shared and learnable linear projection to embed each patch into a hidden space $\mathbf{Z}_{i,(p)}\in\mathbb{R}^{D\times N}$, where $D$ is set to match the dimension used in the subsequent model.
3.2.2 Statistical Embedding
@@
Previous work primarily utilized hard textual prompts to translate dataset descriptions and statistical information [18; 21]. Due to the fact that time series patches and textual prompts come from two different sources, aligning them has proven to be a complex task [12]. In this work, we move away from using hard textual prompts to describe time series statistics. Instead, we generate crucial statistical information separately for both the entire segment and each non-overlapping patch, including the minimum, median, and maximum values (after normalization), as well as trends (calculated as the last observation value of a patch minus its first). This information is represented as $\mathbf{V}_{i}\in\mathbb{R}^{4\times(N+1)}$. Similarly, we employ a shared and learnable linear projection to embed $\mathbf{V}_{i}$ into a hidden space, denoted by $\mathbf{Z}_{i,(v)}\in\mathbb{R}^{D\times(N+1)}$, where $\mathbf{Z}_{i,(v)}=[\mathbf{z}_{i,(v_{g})},\mathbf{Z}_{i,(v_{p})}]$, within this embedding, the first embedding $\mathbf{z}_{i,(v_{g})}\in\mathbb{R}^{D}$ represents the statistical information for the entire time series. The remaining $\mathbf{Z}_{i,(v_{p})}\in\mathbb{R}^{D\times N}$ pertains to each individual patch.
3.2.3 Missing Embedding
@@
Adapting to the missing data pattern is also a crucial aspect. The missing rate of each patch accurately reflects the distribution of missing data across the target time series. Therefore, we have also designed a Missing Embedding $\mathbf{z}_{i,(m)}\in\mathbb{R}^{D\times 1}$. It is a randomly-initialized and learnable parameter. We multiply it with the corresponding patch’s mask ratio $\mathbf{r}_{i}\in\mathbb{R}^{1\times N}$ and add it to the final embedding of the patching information: $\mathbf{E}_{i,(p)}=\mathbf{Z}_{i,(p)}+\mathbf{Z}_{i,(v_{p})}+\mathbf{z}_{i,(m)}\times\mathbf{r}_{i}$.
3.2.4 Domain-Specific Embedding
In some instances, we may need our one-for-all model to function within a specific domain for which we have training data, and we also want to preserve its capability for cross-domain generalization. To achieve this, we explicitly introduce a new embedding, $\mathbf{k}\in\mathbb{R}^{D}$. We anticipate that this embedding will encapsulate unique information from the training data throughout the training process and serve as a repository of domain-specific knowledge. We insert this embedding before the patch embeddings $\mathbf{E}_{i,(p)}$. By integrating the domain-specific embedding and the global statistical embedding with patch embeddings $\mathbf{E}_{i,(p)}$, the final embeddings for input $\mathbf{x}_{i}$ is expressed as $\mathbf{E}_{i}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)}]$, where $\mathbf{E}_{i}\in\mathbb{R}^{(N+2)\times D}$.
3.2.5 Output Layer
@@
After passing through the PLM backbone, we discard the segments corresponding to the two added prefixes: the domain-specific embedding and the variable-specific embedding. This means that we use only the representations of the $N$ patches corresponding to the time series as inputs for the model’s final layer. Although the first two embeddings also participate in the attention calculations, they are not included in the final input. We then flatten the final representation to obtain $\mathbf{h}^{(\text{Layer})}_{i}\in\mathbb{R}^{N\times D}$. Finally, we linearly map these representations back to their original dimensions, obtaining the model outputs $\mathbf{o}_{i}\in\mathbb{R}^{L}$.
3.2.6 One-For-All Model Training
To further enhance our model’s adaptability to different missing patterns, we have incorporated a Contrastive Learning module into our training process. The principle behind this module is straightforward: it ensures that the model can learn relatively similar representations of the same patch, when subjected to various missing rates and patterns. During training, for each input $\mathbf{x}_{i}\in\mathbb{R}^{L}$, we generate two random masks, $\mathbf{m}_{i,1}$ and $\mathbf{m}_{i,2}\in\mathbb{R}^{L}$, to mask the inputs. These two masked time series are then simultaneously used as inputs for the PLM. Each patch outputs a representation and an output head is utilized for contrastive learning. As illustrated in Figure 3, we treat the representations of each patch under different masks as positive samples for each other, while representations from other time series and other patches from the same series under different masks are treated as negative samples. To filter out some similar samples, we do not use representations from other patches of the same time series under the same masks as negative samples. We use the InfoNCE loss [27] combined with the Mean Squared Error (MSE) loss between the model outputs and the original time series as the loss function for the one-for-all model (see Appendix B.3 for details). For training parameters, we freeze the self-attention layers of the PLM and only train the layer normalization, word positional embeddings, and feed-forward neural networks within the PLM.

### v3

@@ Abstract
Time series imputation is critical for many real-world applications and has been widely studied. However, existing models often require specialized designs tailored to specific missing patterns, variables, or domains which limits their generalizability. In addition, current evaluation frameworks primarily focus on domain-specific tasks and often rely on time-wise train/validation/test data splits, which fail to rigorously assess a model’s ability to generalize across unseen variables or domains. In this paper, we present NuwaTS, a novel framework that repurposes Pre-trained Language Models (PLMs) for general time series imputation. Once trained, NuwaTS can be applied to impute missing data across any domain. We introduce specialized embeddings for each sub-series patch, capturing information about the patch, its missing data patterns, and its statistical characteristics. By combining contrastive learning with the imputation task, we train PLMs to create a versatile, one-for-all imputation model. Additionally, we employ a plug-and-play fine-tuning approach, enabling efficient adaptation to domain-specific tasks with minimal adjustments. To evaluate cross-variable and cross-domain generalization, we propose a new benchmarking protocol that partitions the datasets along the variable dimension. Experimental results on over seventeen million time series samples from diverse domains demonstrate that NuwaTS outperforms state-of-the-art domain-specific models across various datasets under the proposed benchmarking protocol. Furthermore, we show that NuwaTS generalizes to other time series tasks, such as forecasting. Our codes are available at https://github.com/Chengyui/NuwaTS.
@@ 1 Introduction
Time series data are pervasive across numerous fields, including transportation (Li et al., 2018), healthcare (Tonekaboni et al., 2020), education (Mao et al., 2024), and meteorology (Bi et al., 2023). However, real-world time series data often suffer from missing values. Incomplete data complicate various time series applications such as forecasting and classification, ultimately degrading the quality of data analysis (Ma et al., 2020). This makes time series imputation especially critical (Wang et al., 2024a). Traditionally, time series imputation methods have leaned heavily on statistical techniques like mean imputation and interpolation (Van Buuren & Groothuis-Oudshoorn, 2011). Yet, with the rise of deep learning, there has been a notable shift towards deep learning models for imputation tasks (Cao et al., 2018; Du et al., 2023).
In traditional deep learning-based imputation models, the typical protocol involves training a model on relatively complete or incomplete time series data within a closed domain from historical records. During the testing phase, the model is validated or tested on newly observed incomplete data, which are future observations of the variables from the training set (Cao et al., 2018; Wu et al., 2022) (Figure 1(b)). However, this approach presents a key limitation: the lack of cross-variable generalization capability. Models trained in this manner may struggle to extend to variables not observed in the training set. For instance, in one particular factory, a newly launched production line generates several new time series data, and it’s uncertain whether the model trained on existing production lines can generalize effectively to those new ones. A more challenging scenario arises when we lack training data from a specific domain and must rely on a model trained from other domains for imputation. This introduces a more complex cross-domain generalization problem (Jin et al., 2022; Wilson et al., 2020). For example, generalizing from traffic time series to factory time series presents significant challenges in transferring learned knowledge across fundamentally different domains (Figure 1(a)).
Figure 1: (a) Cross-variable and cross-domain generalization: time series data across different variables and domains may exhibit both shared and distinct patterns. (b) The conventional train/validation/test division protocol of partitioning datasets along the time dimension. (c) Variable-wise division: the proposed approach trains, validates, and tests models on distinct sets of variables, ensuring the model’s ability to generalize across unseen variables during deployment.
In this paper, we propose a novel framework and benchmarking methodology specifically designed to address and assess time series imputation methods that are capable of both cross-variable and cross-domain generalization. We draw inspiration from the success of foundation models trained on diverse datasets spanning multiple domains (Kirillov et al., 2023; Brown et al., 2020; Peng et al., 2023). These models have demonstrated remarkable ability to generalize across tasks and domains using techniques like fine-tuning (He et al., 2022; Zhuang et al., 2020) or prompting (Jia et al., 2022; Kirillov et al., 2023). In a similar vein, we propose a method that pushes time series imputation beyond the traditional single-domain paradigm, enabling generalization both across variables and across domains.
Specifically, we introduce NuwaTS, a foundation model designed for incomplete time series imputation. Our model segments the time series into patches and employs novel token designs to capture both statistical information and missing data patterns. Additionally, we leverage contrastive learning combined with missing data reconstruction to train an one-for-all imputation model on time series data from diverse domains. Finally, for cases requiring domain-specific adaptation, we have developed a domain-specific prefix embedding and a plug-and-play fine-tuning mechanism. This mechanism introduces continuous prompts at each layer of the frozen pre-trained foundation model without modifying its weights, enabling efficient domain specialization.
Furthermore, we introduce a novel benchmarking paradigm for time series imputation models. Rather than the conventional time-wise train/validation/test partitioning, we partition multivariate time series data along the variable dimension, allocating different variables to the training, validation, and test sets (Figure 1(c)). This approach closely simulates real-world deployment scenarios, where models trained on one domain are required to impute missing data for entirely new variables and domains, effectively assessing the model’s cross-variable and cross-domain generalization capabilities.
Our contributions are summarized as follows:
@@
    We propose a novel and more practically relevant benchmark which divides the multivariate time series data along the variable dimension for time series imputation, which better assesses a model’s ability to generalize to new data.
@@
    We introduce NuwaTS, designed to handle missing data imputation tasks for any incomplete time series. NuwaTS is trained on data from diverse domains and incorporates a light-weight “plug-and-play” fine-tuning technique that requires minimal data and computational resources, making it capable of mending every incomplete time series.
@@
    Under the proposed benchmarking protocol, the one-for-all NuwaTS consistently outperforms domain-specific state-of-the-art methods in imputation tasks across nearly all missing rates. Moreover, fine-tuned NuwaTS can be extended to time series forecasting, where its forecasting results are comparable to or even better than existing domain-specific time series forecasting models.
@@ 2.1 Incomplete Time Series Imputation
Many time series imputation models are tailored to specific missing data patterns and domains, such as randomly missing traffic time series. For instance, matrix factorization models are designed for multivariate time series imputation (Yu et al., 2016). For a dataset from a specific domain with a particular missing rate, it often requires optimizing a specialized model using a low-rank prior as the optimization target for imputation. Recent advancements in deep learning techniques have shown promising results in addressing missing data imputation. Generative models such as Generative Adversarial Networks (GANs) (Luo et al., 2019; Tashiro et al., 2021) and diffusion models are used to learn the underlying distribution of the incomplete time series. Several architectures, such as Recurrent Neural Networks (RNNs) (Cao et al., 2018; Ma et al., 2019; Liu et al., 2019; Tang et al., 2020) and attention mechanisms (Du et al., 2023), are proposed to capture temporal dependencies within incomplete time series. While achieving good performance on narrow domains and missing data patterns, these models lack versatility and generalizability to diverse domains and missing data patterns.
@@ 2.2 Foundation Model for Time Series
In recent years, foundational models for time series analysis have made notable strides, primarily leveraging pre-trained backbones from NLP and aligning modalities to extend their reasoning capabilities to time series tasks. For example, Gruver et al. (Gruver et al., 2024) discovered that by encoding time series as strings of numerical digits, LLMs can perform time series forecasting with zero-shot capabilities. GPT4TS (Zhou et al., 2024) trains time series models using pre-trained GPT weights. UniTime (Liu et al., 2024a) integrates domain-specific instructions into LLMs, enhancing their generalization across domains. Time-LLM (Jin et al., 2024) employs LLMs’ reasoning by framing time series data and statistical information in textual prompts, aligning time patches with word embeddings via cross-attention mechanisms for improved zero-shot learning. Autotimes (Liu et al., 2024d) utilizes precomputed text embedding as positional embeddings and an autoregressive approach for long-term forecasting. TEST (Sun et al., 2023) uses text-prototype-aligned embeddings to enhance LLMs’ reasoning in time series data. $S^{2}$-LLM (Pan et al., 2024) applies seasonal-trend decomposition to time series and uses semantic space-informed prompting to retrieve appropriate prompts from word token embeddings as prefixes. aLLM4TS (Bian et al., 2024) adapts LLMs for time series representation learning by directly converting individual patches into time series sequences. Chronos (Ansari et al., 2024) trains language models from scratch on a large collection of time series data, using scaling and quantization for tokenization. This model demonstrates strong capabilities in zero-shot probabilistic forecasting. These models primarily focus on time series forecasting and do not specifically address missing data issues or the design of embeddings for missing data patterns.
@@ 3.1 Preliminaries and Problem Statement
$m_{t}=\begin{cases}1,&\text{if }x_{t}\text{ is observed}\\ 0,&\text{otherwise}\end{cases}$ (1)
Suppose we have acquired $N$ time series datasets $\mathcal{D}=\{\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{N}\}$ from a diverse set of domains $\{\mathcal{S}^{1},\mathcal{S}^{2},\ldots,\mathcal{S}^{K}\}$, where $\mathbf{x}^{n}\in\mathbb{R}^{T_{n}}$ represents the $n^{th}$ time series with length $T_{n}$. Our objective is to utilize $\mathcal{D}$ to train an imputation model, denoted as $f_{\mathbf{\Phi}}$, characterized by parameters $\mathbf{\Phi}$. For any incomplete time series $\hat{\mathbf{x}}\in\mathbb{R}^{\hat{T}}$ accompanied by any missing pattern $\hat{\mathbf{m}}\in\mathbb{R}^{\hat{T}}$ from any domain $\hat{\mathcal{S}}$, the model $f_{\mathbf{\Phi}}$ can impute the missing values in $\hat{\mathbf{x}}$ as accurately as possible.
@@ 3.2 Model Architecture
Instance Normalization & Patching Time series from different domains can vary in magnitude and distribution. To address these differences, we apply reversible instance normalization (Kim et al., 2021) to each variable before feeding it into the model, with missing values set to zero. To enhance the model’s ability to recognize domain information, we use a patching technique, dividing each time series segment into non-overlapping patches. These patches are then embedded into a hidden space using a shared, learnable linear projection denoted by $\mathbf{Z}_{i,(p)}\in\mathbb{R}^{D\times N}$, enabling more effective modeling of the time series data across domains.
@@
Statistical Embedding. Previous work primarily used hard textual prompts to translate dataset descriptions and statistical information (Liu et al., 2024a; d; Jin et al., 2024), which proved challenging due to the mismatch between time series patches and textual prompts. In this work, we move away from this approach and instead generate statistical information, such as minimum, median, maximum values, and trends, for both the entire variable (denoted by $\mathbf{z}_{i,(v_{g})}\in\mathbb{R}^{D}$) and individual patches (denoted by $\mathbf{Z}_{i,(v_{p})}\in\mathbb{R}^{D\times N}$). This information is embedded into a hidden space using a shared, learnable linear projection, allowing for better alignment with the time series data.
@@
Missing Embedding. Adapting to the missing data pattern is crucial. We design a Missing Embedding, a learnable parameter that captures the missing rate of each patch. This embedding is multiplied with the corresponding patch’s mask ratio and added to the corresponding embed patches: $\mathbf{E}_{i,(p)}=\mathbf{Z}_{i,(p)}+\mathbf{Z}_{i,(v_{p})}+\mathbf{z}_{i,(m)}\times\mathbf{r}_{i}$, allowing the model to better account for missing data across the target time series.
Domain-Specific Embedding. In cases where the model needs to function within a specific domain while preserving cross-domain generalization, we introduce a domain-specific embedding, $\mathbf{k}\in\mathbb{R}^{D}$. This embedding learns to capture domain-specific knowledge during training and is inserted before the patch embeddings. This embedding is beneficial for the proposed fine-tuning process, as discussed in Section 3.3.
The final input embeddings are expressed as $\mathbf{E}_{i}=[\mathbf{k},\mathbf{z}_{i,(v_{g})},\mathbf{E}_{i,(p)}]$, integrating both domain-specific and global statistical information. This design could help model generalize well on new in-domain data or even out-of-domain data.
Contrastive learning. To improve the model’s adaptability to different missing patterns, we incorporate a Contrastive Learning module into the training process. For each input $\mathbf{x}_{i}$, we generate two random masks with different mask ratios and input the masked time series into the PLM. The module ensures the model learns similar representations for the same patch under different masks, treating them as positive samples, while treating representations from other patches and series as negative samples. We use the InfoNCE (Oord et al., 2018) loss combined with MSE loss to optimize the model (details in Appendix B.3).
@@
Output layer. After passing through the PLM backbone, we discard the domain-specific and variable-wise statistical embeddings, retaining only the $N$ patch representations as inputs for the final layer. While the prefixes contribute to causal attention calculations, they are excluded from the final output. The remaining representations are flattened and linearly mapped back to their original dimensions, producing the model outputs $\mathbf{o}_{i}\in\mathbb{R}^{L}$.